OpenStudy (anonymous):

Fan for best answer! What is the value of x? A. 5 B. 5/2 C. 10 D. 2/5

OpenStudy (anonymous):

I need more context to answer that

OpenStudy (idku):



OpenStudy (anonymous):

It is an isosceles triangle which means that there is one right angle and two other angles that equal to 45 degrees with two equal sides

OpenStudy (anonymous):

so how do I figure it out?

OpenStudy (anonymous):

Do you know Pythagorean's theorem?

OpenStudy (anonymous):

not that well

OpenStudy (anonymous):

ok. It only works for right triangles but it's really important. it's a^2 + b^2 = c^2

OpenStudy (anonymous):

where c is your hypotenuse or for your problem it's x

OpenStudy (anonymous):

so if you have a = 5 and b = 5 then its like saying 5^2 + 5^2 = x^2

OpenStudy (anonymous):

so 25 + 25 = 50 = x^2

OpenStudy (anonymous):

find the value of the square root of 50 and that will be your answer

OpenStudy (anonymous):

still a little lost, but I got the 50

OpenStudy (anonymous):

look at the picture that @idku posted. On it you see a, and a as your legs of the triangle right?

OpenStudy (anonymous):

and where x on your picture is, you see a sqrt 2 right?

OpenStudy (anonymous):

thats the same thing as what you should get by the square root of 50. your a value (or 5 for you) multiplied by the square root of 2

OpenStudy (anonymous):

\[\sqrt{50}=\sqrt{25*2}\] do you see how that works?

OpenStudy (anonymous):

so 10

OpenStudy (anonymous):

so then you can take out the 25 because the \[\sqrt{25}=5\] and then you still have the 2 left in the square root so it's \[5\sqrt{2}\]

OpenStudy (anonymous):

i take it that would make your answer B

OpenStudy (anonymous):

yup, thanks

OpenStudy (anonymous):

easier to understand? haha the way i worded it was probably confusing, sorry.